OpenStudy (emilyjones284):

Express r=6(costheta+sintheta) in rectangular coordinates

OpenStudy (anonymous):

without the 6 it is a circle with center (0,0) and radius 1

OpenStudy (emilyjones284):

Yeah I just have no idea how to do it

OpenStudy (anonymous):

i should know this, let me think one first step would be to square both sides because \(r^2=x^2+y^2\)

OpenStudy (anonymous):

we can do this, just about there

OpenStudy (anonymous):

ok got it ready?

OpenStudy (emilyjones284):

yeah hold on im going to do it on a separate piece of paper

OpenStudy (anonymous):

and it is always true that \(x=r\cos(\theta), y=r\sin(\theta)\) or \(\frac{x}{r}=\cos(\theta), \frac{y}{r}=\sin(\theta)\)

OpenStudy (emilyjones284):

so would x=6cos^2theta and y=6sin^2theta

OpenStudy (anonymous):

you have \[r=6(\cos(\theta)+\sin(\theta))\] replace \(\cos(\theta)\) by \(\frac{x}{r}\) and \(\sin(\theta)\) by \(\frac{y}{r}\) to get \[r=\frac{6x}{r}+\frac{6y}{r}\] so far so good?

OpenStudy (anonymous):

to answer your question, no, it is not that

OpenStudy (emilyjones284):

yeah I think I got it so far

OpenStudy (anonymous):

ok to recap \[\frac{x}{r}=\cos(\theta), \frac{y}{r}=\sin(\theta)\] is always true, so we are just going to substitute in \[r=6(\cos(\theta)+\sin(\theta))\] to get \[r=\frac{6x}{r}+\frac{6y}{r}\] only recapping because i remember being confused the first time i saw it ok now multiply both sides by \(r\)

OpenStudy (anonymous):

we get \[r^2=6x+6y\] and since \(r^2=x^2+y^2\) always we have \[x^2+y^2=6x+6y\] and we are either done, or we can write this in the standard form of a circle

OpenStudy (emilyjones284):

Okay I think i get it but I dont know what happened to it being \[r ^{2}=x ^{2}+y ^{2}\] ?

OpenStudy (anonymous):

that is always true

OpenStudy (emilyjones284):

ohhhh never mind I see what you did

OpenStudy (anonymous):

k good

OpenStudy (emilyjones284):

So would that be in rectangular coordinates?

OpenStudy (anonymous):

yes or you can write \[x^2-6x+y^2-6y=0\]and complete the square to get the standard form of the circle

OpenStudy (anonymous):

gotta run good luck

OpenStudy (emilyjones284):

wait so either one of those answers is in rectangular coordinates?

OpenStudy (emilyjones284):

Okay thank you for your help (: